OpenStudy (anonymous):

Evaluate a^2+b a=1/2 b=-1/2

OpenStudy (anonymous):

putting a = 1/2 and b = -1/2 in a^2+b \[(\frac{1}{2})^2+\frac{-1}{2}= \frac{1}{4}-\frac{1}{2} = \frac{-1}{4}\]

OpenStudy (anonymous):

thank you so much!!! :)

OpenStudy (anonymous):

You are welcome :)